At a sale, a suit is being sold for 33% of the regular price. The sale price is $184.80 .
What is the regular price?

Answer by Maths68(1474)

Let
Regular Price = x
33% Discount of x = 33/100 * x
33% Discount of x = 33x/100
Regular price - 33% Discount = Sale price
x-33x/100=184.80
(100x-33x)/100=184.80
67x/100=184.80
67x=18480
x=18480/67
x=275.82

OR

Let
Regular Price = x
Suit has been sold at 33% off the regular price = 100%-33%=67%
Suit has been sold at 67% of the regular price
67% of x = 184.80
67/100 * x = 184.80
67x=184.80*100
67x=18480
x=18480/67
x=275.82


Regular price of the suit = $275.82